X Company currently buys a part from a supplier for $14.09 per unit but is considering making the part itself next year. This year, they purchased 3,400 units of this part.Estimated costs to make the part are:

Per-Unit Total     
Direct materials $3.96    $13,464   
Direct labor 3.80    12,920   
Variable overhead 4.20    14,280   
Fixed overhead 5.30    18,020   
Total $17.26    $58,684   


Of the estimated fixed overhead, $6,667 are common costs that would be allocated to the part; the rest would be additional fixed overhead costs. X Company currently rents unused factory space to a tenant for $2,500; it will have to use this space to make the part.

3. X Company is somewhat uncertain how many units of the part it will need next year. How many units would make the company indifferent between continuing to buy the part and making it?
Answer #1
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars Make Buy Differential
Purchase cost = 3400*14.09            47,906.00         (47,906.00)
Direct materials = 3400*3.96            13,464.00            13,464.00
Direct Labour =3400*3.80            12,920.00            12,920.00
Variable overhead= 3400*4.20            14,280.00            14,280.00
Relevant Fixed overhead=3400*5.30-6667            11,353.00            11,353.00
Loss of rent            (2,500.00)              2,500.00
Total relevant costs            52,017.00            45,406.00              6,611.00
Company should buy as it will save 6,611
Variable cost in making =3.96+3.80+4.20                    11.96
Relevant Fixed overhead in making            11,353.00
Loss of rent              2,500.00
Variable cost of buy                    14.09
3)
Indifference point = (11353+2500)/(14.09-11.96)              6,503.76
Thus answer is 6,504 Units